OpenStudy (anonymous):

Given the function g(x)=-2x+4, find g(-5).

OpenStudy (hoblos):

just replace (x) by (-5) g(-5) = -2(-5) +4 = 10+4 = 14

OpenStudy (anonymous):

g(-5)=-2(-5)+4=10+4=14

OpenStudy (anonymous):

Thx:)

OpenStudy (hoblos):

yw